# Complex Domain Approach for Reversible Data Hiding and Homomorphic Encryption: General Framework and Application to Dispersed Data

### v1

@@ V-B1 Complexity Analysis
VI Security Analysis
This section provides a security analysis for both proposed protocols. The H[i][i]dden-EG protocol, described in Section V-A, and the H[i][i]dden-AggP protocol, detailed in Section V-B), are informally analyzed against common attacks.
Attack Model: The attacker, denoted by $\mathcal{A}$, is assumed to have full control over the communication channel between the Data Collector (DC) and the sensors ($\mathrm{S}_{j}$) and among the sensors themselves. $\mathcal{A}$ can read, modify, delete, insert, and replay messages. However, $\mathcal{A}$ does not possess any private keys (DC’s ElGamal private key $a$, DC’s Paillier private key $\mathsf{K_{priv}}$), nor does $\mathcal{A}$ know the unique symmetric keys $K_{\mathrm{sym},j}$ shared between DC and each $\mathrm{S}_{j}$. Furthermore, $\mathcal{A}$ does not know the shared, time-dependent watermarking keys $w_{k}$.
Both protocols inherently share several key security properties:
Eavesdropping: Confidentiality of transmitted sensitive information (sensor data $d_{j,k}$, watermarks $w_{k}$, challenge factors $\lambda_{k}$) is assured by the use of strong cryptographic primitives. All data is transmitted in ciphertext form. The security relies on the respective underlying hard problems (DLP for ElGamal, Composite Residuosity for Paillier) and the strength of symmetric encryption. $\mathcal{A}$ cannot recover plaintext values from intercepted ciphertexts.
Replay Attack: The primary defense against replay attacks in both protocols is the time-dependent nature of the watermark $w_{k}$. Since each $w_{k}$ is unique to a specific time $t_{k}$, any replayed message from time $t_{j}$ at a current time $t_{k}$ (where $j\neq k$) will result in the decryption yielding $w_{j}$. When the DC checks this extracted watermark against the expected $w_{k}$ for the current time $t_{k}$, it will detect a mismatch, causing the data to be discarded. This mechanism is effective regardless of the specific public-key encryption scheme used.
False Data Injection (FDI): Both protocols mitigate FDI through the watermark verification mechanism. The integrity and authenticity of the data (whether individual or aggregated) fundamentally rely on $\mathcal{A}$ being unable to provide the correct, secretly agreed-upon watermark $w_{k}$ for the specific time $t_{k}$. Any injected false data that does not correspond to the correct $w_{k}$ will result in DC extracting an incorrect watermark, leading to the immediate discarding of the malicious data. Thus, the security against FDI is primarily based on the secrecy and unpredictability of the shared watermark agreement.
While the above defenses are common, each protocol presents specific cryptographic mechanisms and architectural implications that affect its security against certain attacks:
Man-in-the-Middle (MitM) Attack: For both protocols, confidentiality prevents retrieval. Against modification, any unauthorized tampering with ciphertexts by $\mathcal{A}$ (without possessing the necessary private/symmetric keys) will, with overwhelming probability, result in a corrupted value. This leads to an immediate failure in watermark verification. In addition, for each particular protocol:
  • H[i][i]dden-EG: $\mathcal{A}$’s ability to tamper with a single ElGamal ciphertext is directly thwarted by the semantic security of the scheme and the ultimate watermark verification.
  • H[i][i]dden-AggP: The tree aggregation structure introduces intermediate aggregated ciphertexts. $\mathcal{A}$ could potentially tamper with these. However, any such tampering, even if it does not immediately corrupt the ciphertext’s validity, will ultimately be detected by the final watermark verification at the DC, as the homomorphic property only works for valid operations on valid plaintexts. Discarding or delaying messages remains a denial-of-service issue not directly prevented by encryption, though time-dependent watermarks can aid detection.
Masquerading Attack: $\mathcal{A}$ cannot successfully impersonate a sensor $\mathrm{S}_{j}$ without knowing the correct, time-dependent watermark $w_{k}$. This is the primary defense. Then, for each protocol:
  • H[i][i]dden-EG: DC’s public key encryption ensures that only the legitimate DC can decrypt messages. Impersonation by $\mathcal{A}$ requires forging a message that encrypts to the correct watermarked data.
  • H[i][i]dden-AggP: $\mathcal{A}$ also cannot correctly decrypt the symmetrically encrypted $\lambda_{k}$ received from DC without knowing $K_{\mathrm{sym},j}$, which prevents it from correctly participating in the data transformation required for valid messages. The Paillier public key does not provide origin authentication for individual sensor messages, but the watermark provides authentication for the data content itself.
Role of $\lambda_{k}$ (Challenge Factor): In both protocols, the confidentiality of $\lambda_{k}$ (from $\mathcal{A}$) prevents $\mathcal{A}$ from directly influencing or predicting the exact multiplicative factor used in the plaintext domain (in H[i][i]dden-AggP) or encrypted domain (in H[i][i]dden-EG), which, combined with the watermark, makes forgery harder. Its varying nature each round also contributes to freshness.
VII Conclusion
@@
Another future research direction is the potential application of the H[i][i]dden approach to multimedia domains. By adapting integer-based media formats to the complex domain for embedding, this method could offer a novel approach to traditional image, audio, or video watermarking challenges, taking advantage of the inherent mixing properties and high capacity of the H[i][i]dden scheme.
Appendix A Generalized Vector-based Approach for RDH
where a $2\times 2$ matrix makes the appropriate transformation.
Remark 11.
Appendix B Modular Arithmetic over Gaussian Integers
It is worth pointing out that Gaussian prime numbers have some differences compared to the regular integer case. The obvious one is that Gaussian primes can have both real and imaginary parts. The second main difference is that not all integer primes are also Gaussian primes. For example, prime factorizations exists for both 2 and 5 in $\mathbb{Z}[i]$: 2=(1+i)(1i)2=(1+i)(1-i) and 5=(2+i)(2i)5=(2+i)(2-i). The integer primes $p$ that are also Gaussian primes satisfy the following condition $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ as proven, for example, in [7].
@@ Appendix B Modular Arithmetic over Gaussian Integers
It is well known that, in regular integers, for a prime number $p\in\mathbb{Z}$, the elements of the cyclic group $\mathbb{Z}_{p}$ (or $\mathbb{Z}/p\mathbb{Z}$) are $0,1,\dots,p-1$, while the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$ has the same elements, excluding 0. The order of the corresponding cyclic groups $\mathbb{Z}[i]_{\pi}$ and ${\mathbb{Z}[i]}^{*}_{\pi}$ is not so straightforward when $\pi$ is a Gaussian prime. For Gaussian integers, given a prime number $\pi\in\mathbb{Z}[i]$, the norm $\mathrm{N}(\pi)$ determines the order of the additive cyclic group $\mathbb{Z}[i]_{\pi}$, whereas $\mathrm{N}(\pi)-1$ is the order of the multiplicative cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ (because of the exclusion of 0). Proofs for these facts can also be found in [7].
As a practical example, the selection of representatives for the cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ for $\pi=7$, which is both an integer and a Gaussian prime, is discussed below. First, we need to take into account that many possible selections of representatives exists, similar to the integer case. For example, 13 and 6 refer to the same equivalence class in both ${\mathbb{Z}}^{*}_{7}$ and ${\mathbb{Z}[i]}^{*}_{7}$, since 13 and 6 differ in a multiple of 7 (in fact, exactly 7). The only relevant consideration for the selection is not to include more than one representative for the same class. A method for such a selection is described in [7] both when the divider has real and imaginary parts different from zero, and when the imaginary part is zero (the case of the Gaussian prime $\pi=7$).
@@
Appendix C Numerical Example for Watermark Embedding, Extraction and Data Recovery
$\lambda_{1}^{-1}=\frac{3-2i}{3^{2}+2^{2}}=\frac{3}{13}-\frac{2}{13}i.$
The corresponding embedded data are the following:
@@ Appendix C Numerical Example for Watermark Embedding, Extraction and Data Recovery
Remark 13.
At the receiver end, the following computation would be performed:
Thus, taking the real and the imaginary parts after the multiplication recovers both the original data and the watermark, respectively.
$S_{1}=\sum_{j=1}^{3}d_{j,1}=d_{1,1}+d_{1,2}+d_{1,3}=5+8+17=30.$
The receiver end would compute the following:

### v2

@@ 5.3.1 Complexity Analysis
6 Formal Security and Privacy Analysis
In this section, we present to a formal theorem-proof evaluation of the proposed protocols. We demonstrate the resilience of the H[i][i]dden framework against the attack model defined in Section 5.1.3 through rigorous cryptographic proofs.
6.1 Confidentiality and Privacy Proofs
Theorem 1 (Resistance to Eavesdropping).
The H[i][i]dden framework ensures the confidentiality of $d_{k}$ (or $d_{j,k}$), $w_{k}$, and $\lambda_{k}$ against a passive eavesdropper $\mathcal{A}$ under the assumption that the Gaussian ElGamal and Paillier cryptosystems are IND-CPA (Indistinguishability under Chosen Plaintext Attack) secure.
Proof.
Suppose there exists an adversary $\mathcal{A}$ that can recover any part of $d_{k}$ (or $d_{j,k}$), $w_{k}$, or $\lambda_{k}$ from the transmitted ciphertexts with non-negligible advantage. We can construct a simulator $\mathcal{B}$ that uses $\mathcal{A}$ to break the semantic security (IND-CPA) of the underlying cryptosystems.
[…]
In H[i][i]dden-AggP, the individual watermarked data points $\delta_{j,k}^{\prime}$ are protected by the IND-CPA security of the Paillier cryptosystem. The challenge factor $\lambda_{k}$ is transmitted to each sensor $\mathrm{S}_{j}$ using symmetric encryption with a unique key $K_{\mathrm{sym},i}$, whose secrecy is established during initialization via Paillier encryption. If $\mathcal{A}$ could recover $\lambda_{k}$ or any component of the sensor data, the simulator $\mathcal{B}$ would be able to break the underlying semantic security of either the Paillier scheme or the symmetric cipher. Given the assumed security of these primitives, no such adversary $\mathcal{A}$ can exist, and confidentiality is maintained. ∎
Theorem 2 (Aggregation Privacy).
The H[i][i]dden-AggP protocol preserves the privacy of individual sensor readings $d_{j,k}$ against honest-but-curious partial aggregators $\mathrm{S}_{\ell}$.
Proof.
In H[i][i]dden-AggP, a partial aggregator $\mathrm{S}_{\ell}$ receives pairs of ciphertexts $(c_{j,k,R},c_{j,k,I})$, where $c_{j,k,R}=\mathrm{Enc}_{\mathsf{K_{pub}}}(\operatorname{Re}(\delta^{\prime}_{j,k}),r_{j,k,R})$ and $c_{j,k,I}=\mathrm{Enc}_{\mathsf{K_{pub}}}(\operatorname{Im}(\delta^{\prime}_{j,k}),r_{j,k,I})$. The aggregator computes the product component-wise: $C_{\mathrm{sum},R}=\prod_{j}c_{j,k,R}\,(\operatorname{\mathrm{mod}}n^{2})$ and $C_{\mathrm{sum},I}=\prod_{j}c_{j,k,I}\,(\operatorname{\mathrm{mod}}n^{2})$. Due to the IND-CPA security of the Paillier cryptosystem, the aggregator cannot distinguish between the encryption of a valid component (real or imaginary) and an encryption of a random value in $\mathbb{Z}_{n}$. Since $\mathrm{S}_{\ell}$ does not possess the private key $\mathsf{K_{priv}}$, the view of $\mathrm{S}_{\ell}$ consists solely of pairs of ciphertexts which reveal zero information about the underlying component values. Consequently, the aggregator is unable to learn any information about the Gaussian integer $\delta^{\prime}_{j,k}=\lambda_{k}(d_{j,k}+iw_{k})$, ensuring that individual readings $d_{j,k}$ remain private during the aggregation process. ∎
6.2 Integrity, Provenance, and Active Attack Proofs
Theorem 3 (Resistance to Replay Attacks).
An active adversary $\mathcal{A}$ cannot successfully replay a ciphertext (or a component thereof) captured at time $t_{k^{\prime}}$ to be accepted at time $t_{k}$ (where $k\neq k^{\prime}$) with non-negligible probability.
Proof.
At each time sample $t_{k}$, the Data Collector expects a specific Gaussian integer result $\delta^{\prime}_{k}$ (or $\sigma^{\prime}_{k}$) that, upon extraction with the current secret challenge factor $\lambda_{k}$, yields the expected time-dependent watermark $w_{k}$ (or $Nw_{k}$ in the aggregated case). We distinguish two cases:
  1. 1.
    Individual case (H[i][i]dden-EG): If $\mathcal{A}$ replays a ciphertext from $t_{k^{\prime}}$ at time $t_{k}$, the DC recovers $\delta^{\prime}_{k^{\prime}}=\lambda_{k^{\prime}}\delta_{k^{\prime}}$. The extraction yields:
    $\tilde{\delta}_{k}=\frac{\delta^{\prime}_{k^{\prime}}}{\lambda_{k}}=\frac{\lambda_{k^{\prime}}}{\lambda_{k}}(d_{k^{\prime}}+iw_{k^{\prime}}).$
    Since $\lambda_{k}$ and $\lambda_{k^{\prime}}$ are secret and refreshed every round, the resulting ratio $\lambda_{k^{\prime}}/\lambda_{k}$ is unknown to $\mathcal{A}$, making $\operatorname{Im}(\tilde{\delta}_{k})$ effectively random. The condition $\operatorname{Im}(\tilde{\delta}_{k})=w_{k}$ is satisfied only with probability $2^{-B}$.
  2. 2.
    Aggretated case (H[i][i]dden-AggP): If $\mathcal{A}$ captures a sensor’s component ciphertexts $(c_{j,k,R},c_{j,k,I})$ at $t_{k^{\prime}}$ and replays them into the aggregate for $t_{k}$, the decrypted aggregate sum is:
    $\tilde{\sigma}_{k}=\left(\sum_{l\neq j}\lambda_{k}\delta_{l,k}\right)+\lambda_{k^{\prime}}\delta_{j,k^{\prime}}.$
    The DC extracts the aggregate watermark:
    $\operatorname{Im}(\tilde{\sigma}_{k}/\lambda_{k})=(N-1)w_{k}+\operatorname{Im}\left(\frac{\lambda_{k^{\prime}}}{\lambda_{k}}\delta_{j,k^{\prime}}\right).$
    For the aggregate to be accepted, the condition $\operatorname{Im}(\tilde{\sigma_{k}}/\lambda_{k})=Nw_{k}$ must hold, implying that
    $\operatorname{Im}\left(\frac{\lambda_{k^{\prime}}}{\lambda_{k}}(d_{j,k^{\prime}}+iw_{k^{\prime}})\right)=w_{k}.$
    The probability of an accidental collision where the resulting imaginary part matches the current expected watermark $w_{k}$ is exactly $2^{-B}$.
In both cases, the attacker must satisfy a condition involving three independent, secret, and time-varying parameters: the old watermark $w_{k^{\prime}}$, the old challenge $\lambda_{k^{\prime}}$, and the new challenge $\lambda_{k}$. Thus, detection occurs with probability $1-2^{-B}$ in either scenario. ∎
Theorem 4 (Resistance to Man-in-the-Middle (MitM) Attacks).
Any unauthorized modification of a legitimate ciphertext by $\mathcal{A}$ will be detected by the DC with overwhelming probability, preserving data integrity.
Proof.
In the H[i][i]dden framework, the DC chooses a modulus $p$ (for H[i][i]dden-EG) or $n$ (for H[i][i]dden-AggP) sufficiently large to ensure exact recovery of the Gaussian integer value $\delta^{\prime}_{k}=\lambda_{k}\delta_{k}$ or the aggregate $\sigma^{\prime}_{k}=\lambda_{k}\sigma_{k}$, respectively, upon decryption of a valid ciphertext. Consider a MitM adversary $\mathcal{A}$ who intercepts a ciphertext of $\delta^{\prime}_{k}$ (for H[i][i]dden-EG) or $\delta^{\prime}_{j,k}$ (for H[i][i]dden-AggP) and transmits a modified version, resulting in an induced error $\epsilon=e_{R}+ie_{I}$ in the plaintext domain after decryption.
  1. 1.
    Individual case (H[i][i]dden-EG): The DC extracts:
    $\tilde{\delta}_{k}=\frac{\delta^{\prime}_{k}+\epsilon}{\lambda_{k}}=\delta_{k}+\frac{\epsilon}{\lambda_{k}}.$
    Expanding using $\lambda_{k}^{-1}=(a_{k}-ib_{k})/(a_{k}^{2}+b_{k}^{2})$:
    $\tilde{\delta}_{k}=(d_{k}+iw_{k})+\frac{(e_{R}a_{k}+e_{I}b_{k})+i(e_{I}a_{k}-e_{R}b_{k})}{a_{k}^{2}+b_{k}^{2}}.$
    The extracted watermark is (imaginary part):
    $\tilde{w}_{k}=w_{k}+\frac{e_{I}a_{k}-e_{R}b_{k}}{a_{k}^{2}+b_{k}^{2}}.$
  2. 2.
    Aggregated case (H[i][i]dden-AggP): The DC extracts the aggregate
    $\tilde{\sigma}_{k}=\frac{\sigma^{\prime}_{k}+\epsilon}{\lambda_{k}}=\sigma_{k}+\frac{\epsilon}{\lambda_{k}}.$
    Since $\sigma_{k}=S_{k}+iNw_{k}$ (Section 3.3):
    $\tilde{\sigma}_{k}=(S_{k}+iNw_{k})+\frac{(e_{R}a_{k}+e_{I}b_{k})+i(e_{I}a_{k}-e_{R}b_{k})}{a_{k}^{2}+b_{k}^{2}}.$
    The extracted watermark is (imaginary part divided by $N$):
    $\tilde{w}_{k}=w_{k}+\frac{e_{I}a_{k}-e_{R}b_{k}}{N(a_{k}^{2}+b_{k}^{2})}.$
In both cases, for the modification to go undetected, the induced imaginary shift must be zero, implying $e_{I}a_{k}-e_{R}b_{k}=0$. Since $\mathcal{A}$ blindly manipulates ciphertexts without knowledge of $(a_{k},b_{k})$, the attacker cannot choose a modification that purposely aligns $\epsilon$ with the secret “slope” of $\lambda_{k}$. For a large secret space, the probability of an accidental collision is negligible, ensuring protection against MitM modifications. ∎
Theorem 5 (Resistance to Masquerading and False Data Injection (FDI)).
The H[i][i]dden framework ensures resistance to masquerading and FDI attacks; a malicious entity $\mathcal{A}$ cannot successfully inject fraudulent readings $d_{\mathrm{fake}}$ that the DC accepts as originating from a legitimate sensor $\mathrm{S}_{j}$.
Proof.
An FDI attack differs from MitM modification in that $\mathcal{A}$ attempts to inject an entirely synthetic reading $d_{\mathrm{fake}}$ without access to a valid message’s ciphertext. To succeed, $\mathcal{A}$ must generate a ciphertext that decrypts to a specific target value $\delta^{\prime}_{\mathrm{target}}=\lambda_{k}(d_{\mathrm{fake}}+iw_{k})$.
As established in the Attack Model, $\mathcal{A}$ does not possess the secret challenge factor $\lambda_{k}$ nor the current watermark $w_{k}$. In the H[i][i]dden-EG protocol, $\lambda_{k}$ is transmitted to the sensor $\mathrm{S}_{j}$ encrypted using the DC’s public key; its confidentiality is thus guaranteed by the semantic security of the Gaussian ElGamal scheme. Similarly, in the H[i][i]dden-AggP protocol, $\lambda_{k}$ is protected by a unique symmetric key $K_{\mathrm{sym},j}$ shared between the DC and $\mathrm{S}_{j}$.
Consequently, $\mathcal{A}$ is unable to learn $\lambda_{k}$ by eavesdropping in either protocol. Without these secrets, $\mathcal{A}$ would need to guess both the specific multiplicative factor $\lambda_{k}$ and the expected watermark $w_{k}$ for the current time frame. If the secret space for $(a_{k},b_{k})$ has cardinality $|S_{\lambda}|$ and the watermark space has cardinality $2^{B}$, the probability of $\mathcal{A}$ successfully crafting an injected value that survives the DC’s division and watermark extraction test is approximately:
$P(\mathrm{success})\approx\frac{1}{|S_{\lambda}|}\cdot\frac{1}{2^{B}}.$
For standard cryptographic parameters, this probability is negligible. Thus, the successful extraction of $w_{k}$ serves as a proof of origin and authenticity, fulfilling the requirement for data provenance. ∎
Theorem 6 (Perfect Reversibility and Accuracy).
The H[i][i]dden framework provides zero-distortion reconstruction of the host data if the range bounds $\max(|\operatorname{Re}(\delta^{\prime}_{k})|,|\operatorname{Im}(\delta^{\prime}_{k})|)<p/2$ (for H[i][i]dden-EG) and $\max(|\operatorname{Re}(\sigma^{\prime}_{k})|,|\operatorname{Im}(\sigma^{\prime}_{k})|)<n/2$ (for H[i][i]dden-AggP) are satisfied.
Proof.
The correctness of the reconstruction relies on the modular decryption recovering the exact Gaussian integer result of the watermarking process:
  1. 1.
    Individual case (H[i][i]dden-EG): Decryption modulo $p$ in a centered representative system yields the exact value $\delta^{\prime}_{k}=\lambda_{k}(d_{k}+iw_{k})$ if its real and imaginary parts satisfy $|\operatorname{Re}(\delta^{\prime}_{k})|<p/2$ and $|\operatorname{Im}(\delta^{\prime}_{k})|<p/2$.
  2. 2.
    Aggregated case (H[i][i]dden-AggP): Component-wise decryption modulo $n$ recovers the aggregate components exactly if $|\operatorname{Re}(\sigma^{\prime}_{k})|<n/2$ and $|\operatorname{Im}(\sigma^{\prime}_{k})|<n/2$.
In both scenarios, once the exact Gaussian integer ($\delta^{\prime}_{k}$ or $\sigma^{\prime}_{k}$) is recovered, standard complex division by the current secret challenge factor $\lambda_{k}$ in the field $\mathbb{C}$ yields the exact watermarked data $\delta_{k}=d_{k}+iw_{k}$ or the aggregate $\sigma_{k}=S_{k}+iNw_{k}$. The host data ($d_{k}$ or $S_{k}$) is then perfectly reconstructed by taking the real part of the result, ensuring zero distortion. ∎
In short, the formal evaluation conducted through Theorems 16 provides a comprehensive verification of the system requirements and attack model established in Section 5.1. The robustness against eavesdropping (Theorem 1) and the guarantee of aggregation privacy (Theorem 2) satisfy the requirements of confidentiality and privacy. The probabilistic detection of replay (Theorem 3) and the algebraic detection of MitM modifications (Theorem 4) fulfill the integrity requirement. Furthermore, the analysis of masquerading and FDI resistance (Theorem 5) validates the data provenance and authenticity properties by binding successful watermark extraction to the shared secrets. Finally, the accuracy proof (Theorem 6) confirms that the protocols achieve perfect reversibility under defined conditions. Consequently, the H[i][i]dden framework is formally proven to be secure against both passive and active adversarial capabilities $\mathcal{A}$ within the distributed environment.
7 Conclusion
@@
Another future research direction is the potential application of the H[i][i]dden approach to other domains, such as media information or non-dispersed time series data. By adapting integer-based media formats to the complex domain for embedding, this method could offer a novel approach to traditional image, audio, or video watermarking challenges, taking advantage of the inherent mixing properties and high capacity of the H[i][i]dden scheme.
Acknowledgments
This work has been supported, in part, by the “SAFE” (PID2024-156914OB-C41) and “SECURING” (PID2021-125962OB-C31) grants, funded by the Ministry of Science and Innovation thorugh the AEI and the European Regional Development Fund (ERDF), and by the ARTEMISA International Chair of Cybersecurity (C057/23) and the DANGER Strategic Project of Cybersecurity (C062/23), both funded by the Spanish National Institute of Cybersecurity through the European Union—NextGenerationEU and the Recovery, Transformation, and Resilience Plan.
Appendix A Modular Arithmetic over Gaussian Integers
It is worth pointing out that Gaussian prime numbers have some differences compared to the regular integer case. The obvious one is that Gaussian primes can have both real and imaginary parts. The second main difference is that not all integer primes are also Gaussian primes. For example, prime factorizations exists for both 2 and 5 in $\mathbb{Z}[i]$: 2=(1+i)(1i)2=(1+i)(1-i) and 5=(2+i)(2i)5=(2+i)(2-i). The integer primes $p$ that are also Gaussian primes satisfy the following condition $p\equiv 3\,(\operatorname{\mathrm{mod}}4)$ as proven, for example, in (Conrad, 2016).
@@ Appendix A Modular Arithmetic over Gaussian Integers
It is well known that, in regular integers, for a prime number $p\in\mathbb{Z}$, the elements of the cyclic group $\mathbb{Z}_{p}$ (or $\mathbb{Z}/p\mathbb{Z}$) are $0,1,\dots,p-1$, while the multiplicative cyclic group $\mathbb{Z}^{*}_{p}$ has the same elements, excluding 0. The order of the corresponding cyclic groups $\mathbb{Z}[i]_{\pi}$ and ${\mathbb{Z}[i]}^{*}_{\pi}$ is not so straightforward when $\pi$ is a Gaussian prime. For Gaussian integers, given a prime number $\pi\in\mathbb{Z}[i]$, the norm $\mathrm{N}(\pi)$ determines the order of the additive cyclic group $\mathbb{Z}[i]_{\pi}$, whereas $\mathrm{N}(\pi)-1$ is the order of the multiplicative cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ (because of the exclusion of 0). Proofs for these facts can also be found in (Conrad, 2016).
As a practical example, the selection of representatives for the cyclic group ${\mathbb{Z}[i]}^{*}_{\pi}$ for $\pi=7$, which is both an integer and a Gaussian prime, is discussed below. First, we need to take into account that many possible selections of representatives exists, similar to the integer case. For example, 13 and 6 refer to the same equivalence class in both ${\mathbb{Z}}^{*}_{7}$ and ${\mathbb{Z}[i]}^{*}_{7}$, since 13 and 6 differ in a multiple of 7 (in fact, exactly 7). The only relevant consideration for the selection is not to include more than one representative for the same class. A method for such a selection is described in (Conrad, 2016) both when the divider has real and imaginary parts different from zero, and when the imaginary part is zero (the case of the Gaussian prime $\pi=7$).
@@
When the data to be represented is signed (including samples with both positive and negative components), the representation of Fig. 5 is mandatory. In such a case, a simple shift can map every value of Fig. 4 to its corresponding value of Fig. 5 (or conversely). $\triangle$